OpenStudy (goldrush18):

Solve the equation below:

OpenStudy (goldrush18):

\[\log_{4} x=1+\log_{2} 2x, x>0\]

OpenStudy (goldrush18):

i guess we must solve for x

OpenStudy (goldrush18):

@experimentX do u understand?

OpenStudy (experimentx):

yeah ... solve for x.

OpenStudy (experimentx):

use this property of log \[ \log_a b = \frac{\log_kb}{\log_k a}\]

OpenStudy (goldrush18):

i still dont get it sorry

OpenStudy (experimentx):

sorry ... error there \[ \log_4x = \frac{\log_2k}{\log_24} = \frac{\log_2k}{2} \]

OpenStudy (experimentx):

Oops ... error again, that k is supposed to be x

OpenStudy (experimentx):

if you know other few properties of log then it's should be easy for you!!

OpenStudy (experimentx):

\[ \frac{\log_2x}{2} = \log_22 + \log_22x\] \[ \implies \log_2x = 2\log_22 + 2\log_22x \] \[ \implies \log_2x = \log_22^2 + \log_2(2x)^2 \] \[ \implies \log_2x = \log_2(2^2*(2x)^2) \] \[ \implies x = 2^2*(2x)^2 \]

OpenStudy (experimentx):

\[ \implies x = \frac{1}{16} \]

OpenStudy (goldrush18):

@experimentX thanks

OpenStudy (experimentx):

yw ... though not sure about answer!! check on wolframalpha!!

OpenStudy (goldrush18):

its correct

OpenStudy (goldrush18):

because these are past papers they tell us the answer but they dont show us how to get it